A rope pulls a 200 kg sleigh up a slope at incline angle θ=30 deg, through distance
d=20m. The sleigh and its contents have a total mass of 200kg. The snowy slope is so
slippery that we take it to be frictionless. How much work is done by each force
(Normal, gravitational, rope’s) acting on the sleigh?
1
Expert's answer
2020-12-22T05:04:50-0500

Total mass of content is 200 Kg, angle of inclination is 300,distance of travelled = 20 m

Now we can use the below diagram for it



Now tension force= T= mgsin"\theta"

Work done by tension force="mg sin\theta \times s= 200\times 9.8\times sin 30^0\times 20 = 1960\times 0.5\times 20=19600 J"


Gravitation work="mgh= 200\times 9.8\times 20 \times sin 30^0= 19600" J

Normal reaction force = mgcos"\theta= 200\times 9.8\times 0.866=1697.36 N"